OpenStudy (anonymous):

What is the value of a? 5(a-4)-8a=55 @gorv @SubaruWRX @zpupster

OpenStudy (anonymous):

a=-25

OpenStudy (gorv):

5a-20-8a=55 -3a=75 -a=75/3=25 a=-25

OpenStudy (anonymous):

The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle?

OpenStudy (anonymous):

can u guys help me with that one

OpenStudy (gorv):

let width=x length = x+8 perimeter = sum of length of all side x+x+8 +x=x+8=52

OpenStudy (anonymous):

so 52?

OpenStudy (gorv):

x+x+8 +x+x+8=52

OpenStudy (gorv):

sum this up first

OpenStudy (anonymous):

yea idk how

OpenStudy (gorv):

4x+16=52 now subtract 16 from both side 4x=52-16 4x=36 x=36/4

OpenStudy (anonymous):

soit is 36/4

OpenStudy (gorv):

loll 36/4 =9

OpenStudy (anonymous):

omg thank u so much ill become fan :)

OpenStudy (gorv):

lollll